OpenStudy (anonymous):

find the n 6/9 = n/36

OpenStudy (anonymous):

24/36

OpenStudy (idealist10):

6/9=n/36 simplify 6/9 you get 2/3 2/3=n/36 36/3=12 2*12=n n=24 Answer: 24